A sample of 400 MTH409 students showed that the average time they spent studying Mathematics at home is 30 minutes with a standard deviation of 8 minutes. Find 89.9% confidence interval for estimating the average time a student will spend studying Mathematics. Assume that the sample is taken from a normal population.


1
Expert's answer
2021-11-22T15:39:43-0500

"z=\frac{\overline{x}-\mu}{\sigma/\sqrt n}=\frac{30-\mu}{8/20}"

critical values for 89.9% confidence interval:

"z=\pm 1.275"


then:

"-1.275<\frac{30-\mu}{8/20}<1.275"


"-0.51<30-\mu<0.51"


"29.49<\mu<30.51"